OpenStudy (anonymous):

4sqrt{m-1} - sqrt{15m-1}=0

OpenStudy (anonymous):

\[4\sqrt{m-1} -\sqrt{15m-1} =0\]

OpenStudy (agreene):

m=15

OpenStudy (anonymous):

15

jhonyy9 (jhonyy9):

4sqrt(m-1)=sqrt(15m-1) squared both sides 16(m-1)=15m-1 16m-16=15m-1 16m-15m=-1+16 m=15